OpenStudy (anonymous):

determine if convergent and divergent. prove. integral between -infinity and 6 for 10xe^2x dx

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{6} 10xe ^{2x}dx\]

OpenStudy (freckles):

have you tried using integration by parts yet

OpenStudy (anonymous):

i can do an integration by parts. i just don't know the process to arrive to the conclusion of divergence or convergence..... it is from the chapter of improper integrals.

OpenStudy (freckles):

well we can do integration by parts and then apply the limits to find whether or not the integral converges or not

OpenStudy (freckles):

I guess I will assume you did the integration by parts.... \[\lim_{z \rightarrow - \infty }[ 5x e^{2x}-\frac{5}{2}e^{2x}]_{z}^{6}\] plug in upper and lower limit just like you would if you were evaluating a definite integral

OpenStudy (freckles):

then evaluate the limit

OpenStudy (freckles):

if the limit is a number then it converges if the limit is not a number then it diverges

OpenStudy (freckles):

and yes -inf and inf are not numbers

OpenStudy (anonymous):

cool thanks. got it